1) the pitch of a roof is the number of feet the roof rises for each 12 feet horizontally. if a roof has a pitch of 8, what is its slope expressed as a positive number?
Answer by stanbon(75887)

the pitch of a roof is the number of feet the roof rises for each 12 feet horizontally. if a roof has a pitch of 8, what is its slope expressed as a positive number?
------------------------------
slope = verical/horizontal = 8/12 = 2/3
===============
Cheers,
Stan H.